# Quantum speed limits based on Jensen-Shannon and Jeffreys divergences for general physical processes

### v1

@@ Abstract
We discuss quantum speed limits (QSLs) for finite-dimensional quantum systems undergoing a general physical process. These QSLs were obtained using two families of entropic measures, namely the square root of the Jensen-Shannon divergence, which in turn defines a faithful distance of quantum states, and the square root of the quantum Jeffreys divergence. The results apply to both closed and open quantum systems, and are evaluated in terms of the Schatten speed of the evolved state, as well as cost functions that depend on the smallest and largest eigenvalues of both initial and instantaneous states of the quantum system. To illustrate our findings, we focus on the unitary and nonunitary dynamics of mixed single-qubit states. In the first case, we obtain speed limits à la Mandelstam-Tamm that are inversely proportional to the variance of the Hamiltonian driving the evolution. In the second case, we set the nonunitary dynamics to be described by the noisy operations: depolarizing channel, phase damping channel, and generalized amplitude damping channel. We provide analytical results for the two entropic measures, present numerical simulations to support our results on the speed limits, comment on the tightness of the bounds, and provide a comparison with previous QSLs. Our results may find applications in the study of quantum thermodynamics, entropic uncertainty relations, and also complexity of many-body systems.
@@ I Introduction
Entropic measures also have a prominent role in the subject of quantum information processing, especially with regard to the task of distinguishing quantum states. In this setting, quantum relative entropy (QRE) remains a paradigmatic entropic distinguishability measure of quantum states. QRE finds applications in a wide variety of physical scenarios, including the study of entanglement witnesses Piani (2009), quantum coherence measures Streltsov et al. (2017), asymmetry of quantum states Gour et al. (2009), and also entropy production Landi and Paternostro (2021). Recent results address the interplay of the non-monotonic behavior of QRE and information backflow in quantum non-Markovian dynamics Megier et al. (2021); Settimo et al. (2022).
@@
Here we will discuss the role of Jensen-Shannon and Jeffreys divergences in the characterization of speed limits for open quantum systems. To do so, we investigate the rate of change of these quantities and derive a class of speed limits for general physical processes. We specialize these results to the case of unitary dynamics and noisy quantum channels, thus showing that the bounds depend on the smallest and largest eigenvalues of the density matrix. Importantly, our results fit into the Mandelstam-Tamm class of speed limits. In particular, we set the case of initial single-qubit states, and evaluate the bounds for the depolarizing channel, phase damping channel, and generalized amplitude damping channel. We present numerical simulations to support our theoretical predictions.
@@ II Entropic measures
which will be called from now on the quantum Jensen-Shannon distance (QJSD). Importantly, QJSD is: (i) non-negative, ${D_{JS}}(\rho,\varrho)\geq 0$, with ${D_{JS}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${D_{JS}}(\rho,\varrho)={D_{JS}}(\varrho,\rho)$; (iii) and fullfils the triangular inequality, ${D_{JS}}({\rho_{1}},{\rho_{3}})\leq{D_{JS}}({\rho_{1}},{\rho_{2}})+{D_{JS}}({\rho_{2}},{\rho_{3}})$, for all ${\{{\rho_{l}}\}_{l=1,2,3}}\in\mathcal{S}$.
@@ V.2.1 Depolarizing channel
In Fig. 1, we plot the speed limits ${\tau}_{J,JS}^{\text{QSL}}$ [see Figs. 1(a) and 1(c)] and the normalized relative errors $\widetilde{\delta}_{J,JS}({\tau})$ [see Figs. 1(b) and 1(d)], as a function of the mixedness parameter $r$ and the dimensionless parameter ${\Gamma}{\tau}$, for initial single-qubit states evolving under the depolarizing channel. Figures 1(a) and 1(c) show that, respectively, $\tau_{J}^{\text{QSL}}$} [see Eq. (V.2.1)] and $\tau_{JS}^{\text{QSL}}$} [see Eq. (V.2.1)] exhibit similar qualitative behavior, thus being monotonic functions of the parameters $r$ and $\Gamma\tau$. In both cases, for a fixed value $r\in[0,1)$ ($\Gamma\tau\geq 0$), the QSL time ${\tau}_{J,JS}^{\text{QSL}}$ increases monotonically as a function of $\Gamma\tau\geq 0$ ($0\leq r<1$). In particular, for $0\leq r\lesssim 0.2$, one finds that ${\tau}_{J,JS}^{\text{QSL}}$ approaches small values for the overall dynamics. This means that the closer the initial state is to the center of the Bloch sphere, the faster the system evolves towards to the maximally mixed state.
Figures 1(b) and 1(d) show that $\widetilde{\delta}_{J,JS}({\tau})$ presents similar qualitative behavior for both QJPD and QJSD measures. We note that, for a fixed value $\Gamma\tau\geq 0$, then $\widetilde{\delta}_{J,JS}({\tau})$ monotonically decreases as a function of $r$. In particular, for $0.9\lesssim r<1$ and $\Gamma\tau\gtrsim 3$, one finds that the normalized relative errors approaches small values $\widetilde{\delta}_{J,JS}({\tau})\approx 0$. This suggests that the closer the initial state is to a pure state, that is, the further it is from the fixed point of the nonunitary dynamics, the tighter the bounds will be. Otherwise, the bounds are looser for initial states close to the maximally mixed state, regardless of $\Gamma\tau\geq 0$. We note that, for a given fixed value $0\leq r<1$, each relative error approaches a certain value that remains constant at later times of the dynamics.
@@ V.2.2 Phase damping channel
In turn, for the convex combination ${\varpi^{PD}_{t}}=({\rho_{0}}+{\rho^{PD}_{t}})/2$, we find the time-dependent eigenvalues ${\kappa_{\text{max}/\text{min}}}({\varpi^{PD}_{t}})=(1/2)(1\pm{\nu^{PD}_{t}})$, with
@@
Equation (V.2.2) shows that Jeffreys’ relative entropy does not depend on the initial azimuth angle $\phi$}. We note that QJPD vanishes for incoherent probe states with $\theta=0$ that belong to the $z$-axis of the Bloch sphere. On the one hand, for $\Gamma\tau\ll 1$, we find that QJPD becomes ${D_{J}^{2}}({\rho_{0}},{\rho^{PD}_{\tau}})\approx(r/32){(\Gamma\tau)^{2}}{\sin^{2}}(2\theta)\{(1/2)\ln[(1+r)/(1-r)]+r{\tan^{2}}\theta/(1-{r^{2}})\}$, which in turn scales quadratically with the dimensionless parameter $\Gamma\tau$. On the other hand, for $\Gamma\tau\gg 1$, QJPD approaches the asymptotic value ${D_{J}^{2}}({\rho_{0}},{\rho^{PD}_{\tau}})\approx(r/4){\sin^{2}}{\theta}\ln[(1+r)/(1-r)]$, which is time-independent.
Figure 4: (Color online) Density plot of the normalized relative error ${\widetilde{\delta}_{J}}({\tau})$, as a function of the dimensionless parameters $\Gamma\tau$ and $r\in[0,1)$, for initial single-qubit states subject to the generalized amplitude damping channel [see Sec. V.2.3]. Here, we set $\alpha=\{0,0.1,1\}$ and $\theta=\{0,\pi/4,\pi/2$}, for all $\phi\in[0,2\pi)$.
@@
In Fig. 2, we show the density plots of the entropic QSL time $\tau^{\text{QSL}}_{J,JS}$ [see Figs. 2(a) and 2(c)], and the normalized relative error ${\widetilde{\delta}_{J,JS}}(\tau)$ [see Figs. 2(b) and 2(d)], as a function of the dimensionless parameters $\Gamma\tau$ and $r$, for single-qubit states evolving under the phase damping channel. Here we set initial states with $\theta=\pi/2$ lying in the equatorial plane of the Bloch sphere. Figures 2(a) and 2(c) show that $\tau^{\text{QSL}}_{J}$ and $\tau^{\text{QSL}}_{JS}$ exhibit the same qualitative behavior, monotonically increasing as a function of $\Gamma\tau$ and $r$. It is worth noting that $\tau_{J,JS}^{\text{QSL}}$} approaches small values at earlier times of the dynamics ($0\leq\Gamma\tau\lesssim 2$), regardless the purity of the initial state. In addition, for all $\Gamma\tau\geq 0$, the less pure the probe state, the smaller the entropic speed limit time.
In Figs. 2(b) and 2(d), we plot the normalized relative errors ${\widetilde{\delta}_{J,JS}}(\tau)$ as a function of the dimensionless parameters $\Gamma\tau$ and $r$. On the one hand, for fixed $\Gamma\tau\geq 0$, one finds that ${\widetilde{\delta}_{J,JS}}(\tau)$ decreases as a function of the mixedness parameter $r\in[0,1)$. On the other hand, for fixed $r\gtrsim 0.2$, we have that ${\widetilde{\delta}_{J,JS}}(\tau)$ monotonically decreases as a function of $\Gamma\tau$, and approaches a constant value at later times of the dynamics. The closer the probe state is to the maximally mixed state, the looser the bounds will be, for all $\Gamma\tau\geq 0$. We note that ${\widetilde{\delta}_{J,JS}}(\tau)$ approaches small values at later times of the dynamics for initial states with purity close to the unity.
@@
Figure 6: (Color online) Density plot of the normalized relative error ${\widetilde{\delta}_{JS}}({\tau})$, as a function of the dimensionless parameters $\Gamma\tau$ and $r\in[0,1)$, for initial single-qubit states subject to the generalized amplitude damping channel [see Sec. V.2.3]. Here, we set $\alpha=\{0,0.1,1\}$ and $\theta=\{0,\pi/4,\pi/2$}, for all $\phi\in[0,2\pi)$.
@@ V.2.3 Generalized amplitude damping channel
In this setting, the evolved state becomes ${\rho^{GAD}_{t}}={\sum_{j=0}^{3}}\,{K_{j}}{\rho_{0}}{K_{j}^{\dagger}}=(1/2)\left(\mathbb{I}+{\vec{r}^{\,GAD}_{t}}\cdotp\vec{\sigma}\right)$, with ${\vec{r}^{\,GAD}_{t}}={\eta_{t}^{GAD}}\vec{r}+{\vec{\kappa}^{\,GAD}_{t}}$, where ${\eta_{t}^{GAD}}=\text{diag}(\sqrt{1-{\lambda_{t}}},\sqrt{1-{\lambda_{t}}},1-{\lambda_{t}})$ is a diagonal matrix, while ${\vec{\kappa}^{\,GAD}_{t}}=(0,0,(2\alpha-1){\lambda_{t}})$ is a real-valued tridimensional vector. In particular, for $\Gamma t\gg 1$, one finds the asymptotic vector ${\vec{r}^{\,GAD}_{\infty}}\approx(0,0,2\alpha-1)$, which in turn describes the stationary, mixed, single-qubit state ${\rho^{GAD}_{\infty}}\approx\alpha|0\rangle\langle{0}|+(1-\alpha)|1\rangle\langle{1}|$. This means that the GAD channel shrinks the Bloch sphere along the $z$-axis towards to a fixed point located at $z=2\alpha-1$ Nielsen and Chuang (2000). The instantaneous state exhibits the eigenvalues ${\kappa_{\text{max}/\text{min}}}({\rho^{GAD}_{t}})=(1\pm{r^{GAD}_{t}})/2$, with
@@
In particular, for $t=0$, Eq. (V.2.3) reduces to ${\nu^{GAD}_{0}}=r$, which agrees with the fact that ${\varpi_{0}^{GAD}}={\rho_{0}}$ recovers the initial state. Furthermore, for $\Gamma{t}\gg 1$, we have that ${\nu^{GAD}_{\infty}}\approx(1/2)\sqrt{{(r\cos\theta+2\alpha-1)^{2}}+{r^{2}}{\sin^{2}}\theta}$, thus indicating that the states $\rho_{0}$ and $\rho_{\infty}$ do not commute with each other.
@@
It is noteworthy that, for all $0\leq r<1$ and $\Gamma\tau>0$, we have that ${D_{J,JS}}({\rho_{0}},{\rho^{GAD}_{\tau}})$ remains invariant under the mapping $(\theta,\alpha)\rightarrow(\pi\pm\theta,1-\alpha)$, with $\theta\in[0,\pi]$ and $0\leq\alpha\leq 1$.
The entropic QSL times $\tau_{J,JS}^{\text{QSL}}$} related to the QJPD and QJSD are evaluated through Eqs. (26) and (27), respectively. The expressions are too long to be reported here. In Table 1, we show some useful quantities needed to compute $\tau_{J,JS}^{\text{QSL}}$} for initial single-qubit states evolving under the generalized amplitude damping channel. We remind that the cost functions ${f_{J,JS}}({\rho_{0}},{{\rho^{GAD}_{t}}})$ are given in Eqs. (8) and (9), respectively. It turns out that $\tau_{J,JS}^{\text{QSL}}$} and ${\widetilde{\delta}_{J,JS}}({\tau})$ are symmetric under the mapping $(\theta,\alpha)\rightarrow(\pi\pm\theta,1-\alpha)$, this being an inherited property from QJPD and QJSD. In other words, given two probe states with $(r,{\theta_{1}},\phi,{\alpha_{1}})$ and $(r,{\theta_{2}},\phi,{\alpha_{2}})$, one expects to obtain the same QSL time whenever we have $|{\theta_{1(2)}}-{\theta_{2(1)}}|=\pi$ and ${\alpha_{1}}+{\alpha_{2}}=1$, for all $\Gamma{\tau}>0$.
@@
In Fig. 3, we show the density plots of the entropic QSL time $\tau_{J}^{\text{QSL}}$}, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. We note that, for a fixed value $r\in[0,1)$ of the mixedness parameter, with $\alpha=0.1$ and $\theta=\{0,\pi/4,\pi/2\}$, then $\tau_{J}^{\text{QSL}}$} monotonically increases as a function of the dimensionless parameter $\Gamma\tau\geq 0$ [see Figs. 3(b), 3(e), and 3(h)]. In particular, for $\alpha=0$ and $\alpha=1$, with $\theta=\{\pi/4,\pi/2\}$, we have that the entropic QSL approaches small values at earlier ($\Gamma\tau\ll 1$) and later ($\Gamma\tau\gg 1$) times of the dynamics for initial mixed states, regardless $0<r<1$ [see Figs. 3(d), 3(f), 3(g), and 3(i)]. In this setting, one finds that $\tau_{J}^{\text{QSL}}$} exhibits a non-monotonic behavior as a function of $\Gamma\tau$ along the nonunitary evolution of the two-level system. Indeed, Figs. 3(d), 3(f), 3(g), and 3(i) show that, for fixed value $r\in[0,1)$, the QSL time starts growing but smoothly approaches small values as a function of $\Gamma\tau\geq 0$. It is noteworthy that, for $\alpha=\{0,1\}$ and $\theta=\pi/2$, $\tau_{J}^{\text{QSL}}$} [see Eq. (26)] exhibits the same quantitative behavior, which can be numerically verified in Figs. 3(g) and 3(i).
@@
In Fig. 4, we show the density plots of the normalized relative error ${\widetilde{\delta}_{J}}({\tau})$, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. The entropic QSL bound is loose at earlier times of the dynamics, regardless of the initial state and the value of the parameter $\alpha\in[0,1]$. In addition, Figs. 4(c) and 4(f) show that the bound is also loose for probe states close to the incoherent state $|0\rangle\langle{1}|$, for all $\Gamma\tau\geq 0$. Overall, for $\alpha\ll 1$ and $\theta=\{0,\pi/4,\pi/2\}$, one finds that ${\widetilde{\delta}_{J}}({\tau})$ approaches small values for all $\Gamma\tau\gg 1$ [see Figs. 4(a), 4(b), 4(d), 4(e), 4(g), and 4(h)]. In particular, for $\theta=\pi/2$, tighter entropic QSLs are found at later times of the dynamics whenever the initial state is close to the maximally mixed state, regardless the value of $\alpha=\{0,0.1,1\}$ [see Figs. 4(g), 4(h), and 4(i)].

### v2

@@ Abstract
We discuss quantum speed limits (QSLs) for finite-dimensional quantum systems undergoing general physical processes. These QSLs were obtained using two families of entropic measures, namely the square root of the Jensen-Shannon divergence, which in turn defines a faithful distance of quantum states, and the square root of the quantum Jeffreys divergence. The results apply to both closed and open quantum systems, and are evaluated in terms of the Schatten speed of the evolved state, as well as cost functions that depend on the smallest and largest eigenvalues of both initial and instantaneous states of the quantum system. To illustrate our findings, we focus on the unitary and nonunitary dynamics of mixed single-qubit states. In the first case, we obtain speed limits à la Mandelstam-Tamm that are inversely proportional to the variance of the Hamiltonian driving the evolution. In the second case, we set the nonunitary dynamics to be described by the noisy operations: depolarizing channel, phase damping channel, and generalized amplitude damping channel. We provide analytical results for the two entropic measures, present numerical simulations to support our results on the speed limits, comment on the tightness of the bounds, and provide a comparison with previous QSLs. Our results may find applications in the study of quantum thermodynamics, entropic uncertainty relations, and also complexity of many-body systems.
@@ I Introduction
Entropic measures also have a prominent role in the subject of quantum information processing, especially with regard to the task of distinguishing quantum states. The nonuniqueness of these quantifiers manifests itself in the existence of a zoo of entropies Faist . In this setting, quantum relative entropy (QRE) remains a paradigmatic entropic distinguishability measure of quantum states. QRE finds applications in a wide variety of physical scenarios, including the study of entanglement witnesses Piani (2009), quantum coherence measures Streltsov et al. (2017), asymmetry of quantum states Gour et al. (2009), and also entropy production Landi and Paternostro (2021). Recent results address the interplay of the non-monotonic behavior of QRE and information backflow in quantum non-Markovian dynamics Megier et al. (2021); Settimo et al. (2022).
@@
Here, we discuss the role of Jensen-Shannon and Jeffreys divergences in the characterization of quantum speed limits. To do so, we investigate the rate of change of these quantities and derive a class of speed limits for general physical processes. We specialize these results to the case of unitary dynamics and noisy quantum channels, thus showing that the bounds depend on the smallest and largest eigenvalues of the density matrix. Importantly, our results fit into the Mandelstam-Tamm class of speed limits. In particular, we set the case of initial single-qubit states, and evaluate the bounds for the depolarizing channel, phase damping channel, and generalized amplitude damping channel. We present numerical simulations to support our theoretical predictions.
@@ II Entropic measures
Figure 1: (Color online) Overview of the role played by quantum Jeffreys pseudo-distance (QJPD), ${D_{J}}({\rho},{\varrho})$ [see Eq. (II)], and quantum Jensen-Shannon distance (QJSD), ${D_{JS}}({\rho},{\varrho})$ [see Eq. (II)]. QJPD (blue solid lines) and QJSD (solid red lines) map density matrices of the convex manifold of quantum states $\mathcal{S}$ (gray smooth surface) to nonnegative numbers on the real line $\mathbb{R}$ (black solid line), i.e., ${D_{J,JS}}:({\rho},{\varrho})\in\mathcal{S}\mapsto\mathbb{R}$. Noteworty, QJSD takes into account the “average state” $\varpi$. QJSD is a faithful metric of quantum states, thus linking the distinguishability between $\rho$ and $\varrho$ to the length of a given path that connects these states (black dash-dotted line). QJPD, however, is not a true distance, but serves as a useful entropic measure.
@@
which will be called from now on the quantum Jensen-Shannon distance (QJSD). Importantly, QJSD is: (i) non-negative, ${D_{JS}}(\rho,\varrho)\geq 0$, with ${D_{JS}}(\rho,\varrho)=0$ if and only if $\rho=\varrho$; (ii) symmetric, ${D_{JS}}(\rho,\varrho)={D_{JS}}(\varrho,\rho)$; (iii) and fullfils the triangular inequality, ${D_{JS}}({\rho_{1}},{\rho_{3}})\leq{D_{JS}}({\rho_{1}},{\rho_{2}})+{D_{JS}}({\rho_{2}},{\rho_{3}})$, for all ${\{{\rho_{l}}\}_{l=1,2,3}}\in\mathcal{S}$. In Fig. 1, we show a schematic scenario of both QJPD and QJSD mappings.
@@ V.2.1 Depolarizing channel
In Fig. 2, we plot the speed limits ${\tau}_{J,JS}^{\text{QSL}}$ [see Figs. 2(a) and 2(c)] and the normalized relative errors $\widetilde{\delta}_{J,JS}({\tau})$ [see Figs. 2(b) and 2(d)], as a function of the mixedness parameter $r$ and the dimensionless parameter ${\Gamma}{\tau}$, for initial single-qubit states evolving under the depolarizing channel. Figures 2(a) and 2(c) show that, respectively, $\tau_{J}^{\text{QSL}}$} [see Eq. (V.2.1)] and $\tau_{JS}^{\text{QSL}}$} [see Eq. (V.2.1)] exhibit similar qualitative behavior, thus being monotonic functions of the parameters $r$ and $\Gamma\tau$. In both cases, for a fixed value $r\in[0,1)$ ($\Gamma\tau\geq 0$), the QSL time ${\tau}_{J,JS}^{\text{QSL}}$ increases monotonically as a function of $\Gamma\tau\geq 0$ ($0\leq r<1$). In particular, for $0\leq r\lesssim 0.2$, one finds that ${\tau}_{J,JS}^{\text{QSL}}$ approaches small values for the overall dynamics. This means that the closer the initial state is to the center of the Bloch sphere, the faster the system evolves towards to the maximally mixed state.
Figures 2(b) and 2(d) show that $\widetilde{\delta}_{J,JS}({\tau})$ presents similar qualitative behavior for both QJPD and QJSD measures. We note that, for a fixed value $\Gamma\tau\geq 0$, then $\widetilde{\delta}_{J,JS}({\tau})$ monotonically decreases as a function of $r$. In particular, for $0.9\lesssim r<1$ and $\Gamma\tau\gtrsim 3$, one finds that the normalized relative errors approaches small values $\widetilde{\delta}_{J,JS}({\tau})\approx 0$. This suggests that the closer the initial state is to a pure state, that is, the further it is from the fixed point of the nonunitary dynamics, the tighter the bounds will be. Otherwise, the bounds are looser for initial states close to the maximally mixed state, regardless of $\Gamma\tau\geq 0$. We note that, for a given fixed value $0\leq r<1$, each relative error approaches a certain value that remains constant at later times of the dynamics.
@@ V.2.2 Phase damping channel
In turn, for the convex combination ${\varpi^{PD}_{t}}=(1/2)({\rho_{0}}+{\rho^{PD}_{t}})$, we find the time-dependent eigenvalues ${\kappa_{\text{max}/\text{min}}}({\varpi^{PD}_{t}})=(1/2)(1\pm{\nu^{PD}_{t}})$, with
@@
Equation (V.2.2) shows that the quantum Jeffreys pseudo-distance does not depend on the initial azimuth angle $\phi$}. We note that QJPD vanishes for incoherent probe states with $\theta=0$ that belong to the $z$-axis of the Bloch sphere. On the one hand, for $\Gamma\tau\ll 1$, we find that QJPD becomes ${D_{J}^{2}}({\rho_{0}},{\rho^{PD}_{\tau}})\approx(r/32){(\Gamma\tau)^{2}}{\sin^{2}}(2\theta)\{(1/2)\ln[(1+r)/(1-r)]+r{\tan^{2}}\theta/(1-{r^{2}})\}$, which in turn scales quadratically with the dimensionless parameter $\Gamma\tau$. On the other hand, for $\Gamma\tau\gg 1$, QJPD approaches the asymptotic value ${D_{J}^{2}}({\rho_{0}},{\rho^{PD}_{\tau}})\approx(r/4){\sin^{2}}{\theta}\ln[(1+r)/(1-r)]$, which is time-independent.
Figure 5: (Color online) Density plot of the normalized relative error ${\widetilde{\delta}_{J}}({\tau})$ [see Eq. (33)], as a function of the dimensionless parameters $\Gamma\tau$ and $r\in[0,1)$, for initial single-qubit states subject to the generalized amplitude damping channel [see Sec. V.2.3]. Here, we set $\alpha=\{0,0.1,1\}$ and $\theta=\{0,\pi/4,\pi/2$}, for all $\phi\in[0,2\pi)$.
@@
In Fig. 3, we show the density plots of the entropic QSL time $\tau^{\text{QSL}}_{J,JS}$ [see Figs. 3(a) and 3(c)], and the normalized relative error ${\widetilde{\delta}_{J,JS}}(\tau)$ [see Figs. 3(b) and 3(d)], as a function of the dimensionless parameters $\Gamma\tau$ and $r$, for single-qubit states evolving under the phase damping channel. Here we set initial states with $\theta=\pi/2$ lying in the equatorial plane of the Bloch sphere. Figures 3(a) and 3(c) show that $\tau^{\text{QSL}}_{J}$ and $\tau^{\text{QSL}}_{JS}$ exhibit the same qualitative behavior, monotonically increasing as a function of $\Gamma\tau$ and $r$. It is worth noting that $\tau_{J,JS}^{\text{QSL}}$} approaches small values at earlier times of the dynamics ($0\leq\Gamma\tau\lesssim 2$), regardless the purity of the initial state. In addition, for all $\Gamma\tau\geq 0$, the less pure the probe state, the smaller the entropic speed limit time.
In Figs. 3(b) and 3(d), we plot the normalized relative errors ${\widetilde{\delta}_{J,JS}}(\tau)$ as a function of the dimensionless parameters $\Gamma\tau$ and $r$. On the one hand, for fixed $\Gamma\tau\geq 0$, one finds that ${\widetilde{\delta}_{J,JS}}(\tau)$ decreases as a function of the mixedness parameter $r\in[0,1)$. On the other hand, for fixed $r\gtrsim 0.2$, we have that ${\widetilde{\delta}_{J,JS}}(\tau)$ monotonically decreases as a function of $\Gamma\tau$, and approaches a constant value at later times of the dynamics. The closer the probe state is to the maximally mixed state, the looser the bounds will be, for all $\Gamma\tau\geq 0$. We note that ${\widetilde{\delta}_{J,JS}}(\tau)$ approaches small values at later times of the dynamics for initial states with purity close to the unity.
@@
Figure 7: (Color online) Density plot of the normalized relative error ${\widetilde{\delta}_{JS}}({\tau})$ [see Eq. (33)], as a function of the dimensionless parameters $\Gamma\tau$ and $r\in[0,1)$, for initial single-qubit states subject to the generalized amplitude damping channel [see Sec. V.2.3]. Here, we set $\alpha=\{0,0.1,1\}$ and $\theta=\{0,\pi/4,\pi/2$}, for all $\phi\in[0,2\pi)$.
@@ V.2.3 Generalized amplitude damping channel
In this setting, the evolved state becomes ${\rho^{GAD}_{t}}={\sum_{j=0}^{3}}\,{K_{j}}{\rho_{0}}{K_{j}^{\dagger}}=(1/2)\left(\mathbb{I}+{\vec{r}^{\,GAD}_{t}}\cdotp\vec{\sigma}\right)$, with ${\vec{r}^{\,GAD}_{t}}={\eta_{t}^{GAD}}\vec{r}+{\vec{\kappa}^{\,GAD}_{t}}$, where ${\eta_{t}^{GAD}}=\text{diag}(\sqrt{1-{\lambda_{t}}},\sqrt{1-{\lambda_{t}}},1-{\lambda_{t}})$ is a diagonal matrix, while ${\vec{\kappa}^{\,GAD}_{t}}=(0,0,(2\alpha-1){\lambda_{t}})$ is a real-valued tridimensional vector. In particular, for $\Gamma t\gg 1$, one finds the asymptotic vector ${\vec{r}^{\,GAD}_{\infty}}\approx(0,0,2\alpha-1)$, which in turn describes the stationary, mixed, single-qubit state ${\rho^{GAD}_{\infty}}\approx\alpha|0\rangle\langle{0}|+(1-\alpha)|1\rangle\langle{1}|$. This means that the GAD channel shrinks the Bloch sphere along the $z$-axis towards to a fixed point located at $z=2\alpha-1$ Nielsen and Chuang (2000).
The instantaneous state exhibits the eigenvalues ${\kappa_{\text{max}/\text{min}}}({\rho^{GAD}_{t}})=(1/2)(1\pm{r^{GAD}_{t}})$, with
@@
In particular, for $t=0$, Eq. (V.2.3) reduces to ${\nu^{GAD}_{0}}=r$, which agrees with the fact that ${\varpi_{0}^{GAD}}={\rho_{0}}$ recovers the initial state. Furthermore, for $\Gamma{t}\gg 1$, we have that ${\nu^{GAD}_{\infty}}\approx(1/2)\sqrt{{(r\cos\theta+2\alpha-1)^{2}}+{r^{2}}{\sin^{2}}\theta}$. This indicates that the states $\rho_{0}$ and $\rho_{\infty}$ do not commute with each other, and the asymptotic eigenvalues ${\kappa_{\text{max}/\text{min}}}({\varpi^{GAD}_{\infty}})$ still depend on the parameters $(r,\theta,\alpha)$.
@@
It is noteworthy that, for all $0\leq r<1$ and $\Gamma\tau>0$, we have that ${D_{J,JS}}({\rho_{0}},{\rho^{GAD}_{\tau}})$ remains invariant under the mapping $(\theta,\alpha)\rightarrow(\theta\pm\pi,1-\alpha)$, with $\theta\in[0,\pi]$ and $0\leq\alpha\leq 1$. This property is inherited from Eqs. (V.2.3) and (V.2.3), which means that the eigenvalues ${\kappa_{\text{max}/\text{min}}}({\rho^{GAD}_{t}})$ and ${\kappa_{\text{max}/\text{min}}}({\varpi^{GAD}_{t}})$ exhibit such a symmetry.
The entropic QSL times $\tau_{J,JS}^{\text{QSL}}$} related to the QJPD and QJSD are evaluated through Eqs. (26) and (27), respectively. The expressions are too long to be reported here. In Table 1, we show some useful quantities needed to compute $\tau_{J,JS}^{\text{QSL}}$} for initial single-qubit states evolving under the generalized amplitude damping channel. We remind that the cost functions ${f_{J,JS}}({\rho_{0}},{{\rho^{GAD}_{t}}})$ are given in Eqs. (8) and (9), respectively. It turns out that $\tau_{J,JS}^{\text{QSL}}$} and ${\widetilde{\delta}_{J,JS}}({\tau})$ are symmetric under the mapping $(\theta,\alpha)\rightarrow(\theta\pm\pi,1-\alpha)$, this being an inherited property from QJPD and QJSD. In other words, given two probe states with $(r,{\theta_{1}},\phi,{\alpha_{1}})$ and $(r,{\theta_{2}},\phi,{\alpha_{2}})$, one expects to obtain the same QSL time whenever we have $|{\theta_{1(2)}}-{\theta_{2(1)}}|=\pi$ and ${\alpha_{1}}+{\alpha_{2}}=1$, for all $\Gamma{\tau}>0$.
@@
In Fig. 4, we show the density plots of the entropic QSL time $\tau_{J}^{\text{QSL}}$}, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. We note that, for a fixed value $r\in[0,1)$ of the mixedness parameter, with $\alpha=0.1$ and $\theta=\{0,\pi/4,\pi/2\}$, then $\tau_{J}^{\text{QSL}}$} monotonically increases as a function of the dimensionless parameter $\Gamma\tau\geq 0$ [see Figs. 4(b), 4(e), and 4(h)]. In particular, for $\alpha=0$ and $\alpha=1$, with $\theta=\{\pi/4,\pi/2\}$, we have that the entropic QSL approaches small values at earlier ($\Gamma\tau\ll 1$) and later ($\Gamma\tau\gg 1$) times of the dynamics for initial mixed states, regardless $0<r<1$ [see Figs. 4(d), 4(f), 4(g), and 4(i)]. In this setting, one finds that $\tau_{J}^{\text{QSL}}$} exhibits a non-monotonic behavior as a function of $\Gamma\tau$ along the nonunitary evolution of the two-level system. Indeed, Figs. 4(d), 4(f), 4(g), and 4(i) show that, for fixed value $r\in[0,1)$, the QSL time starts growing but smoothly approaches small values as a function of $\Gamma\tau\geq 0$. It is noteworthy that, for $\alpha=\{0,1\}$ and $\theta=\pi/2$, $\tau_{J}^{\text{QSL}}$} [see Eq. (26)] exhibits the same quantitative behavior, which can be numerically verified in Figs. 4(g) and 4(i).
@@
In Fig. 5, we show the density plots of the normalized relative error ${\widetilde{\delta}_{J}}({\tau})$, as a function of the dimensionless parameters $\Gamma\tau$ and $r$. The entropic QSL bound is loose at earlier times of the dynamics, regardless of the initial state and the value of the parameter $\alpha\in[0,1]$. In addition, Figs. 5(c) and 5(f) show that the bound is also loose for probe states close to the incoherent state $|0\rangle\langle{1}|$, for all $\Gamma\tau\geq 0$. Overall, for $\alpha\ll 1$ and $\theta=\{0,\pi/4,\pi/2\}$, one finds that ${\widetilde{\delta}_{J}}({\tau})$ approaches small values for all $\Gamma\tau\gg 1$ [see Figs. 5(a), 5(b), 5(d), 5(e), 5(g), and 5(h)]. In particular, for $\theta=\pi/2$, tighter entropic QSLs are found at later times of the dynamics whenever the initial state is close to the maximally mixed state, regardless the value of $\alpha=\{0,0.1,1\}$ [see Figs. 5(g), 5(h), and 5(i)].